# AfriMed-QA: A Pan-African, Multi-Specialty, Medical Question-Answering Benchmark Dataset

### v2

@@ Abstract
Recent advancements in large language model (LLM) performance on medical multiple-choice question (MCQ) benchmarks have stimulated interest from healthcare providers and patients globally. Particularly in low-and-middle-income countries (LMICs) facing acute physician shortages and lack of specialists, LLMs offer a potentially scalable pathway to enhance healthcare access and reduce costs. However, their effectiveness in the Global South, especially across the African continent, remains to be established. In this work, we introduce AfriMed-QA 22footnotetext: Our dataset is available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
, the first large-scale Pan-African English multi-specialty medical Question-Answering (QA) dataset, 15,000 questions (open and closed-ended) sourced from over 60 medical schools across 16 countries, covering 32 medical specialties. We further evaluate 30 LLMs across multiple axes including correctness and demographic bias. Our findings show significant performance variation across specialties and geographies, MCQ performance clearly lags USMLE (MedQA). We find that biomedical LLMs underperform general models and smaller edge-friendly LLMs struggle to achieve a passing score. Interestingly, human evaluations show a consistent consumer preference for LLM answers and explanations when compared with clinician answers.
@@ 1 Introduction
To address this gap, we introduce AfriMed-QA, a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias. We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches. While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs.
@@ 2.2 Evaluating LLMs for Health-Specific Tasks
Reddy et al. Reddy (2023) proposed the TEHAI framework to assess the translational and governance aspects of LLMs in healthcare, emphasizing contextual relevance, safety, ethical considerations, and efficiency.
Our evaluation extends TEHAI and work from Singhal et al. Singhal et al. (2022b), and Pfohl et al. Pfohl et al. (2024) by incorporating specialty and region-specific dimensions, ensuring a comprehensive assessment of LLMs. Metrics such as local expertise, harmfulness, and bias are included to cover ethical dimensions and governance. By addressing correctness, omission, hallucination, and reasonableness, we thoroughly evaluate AI systems in healthcare.
@@ 4.2 Quantitative Evaluation
Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore
@@ 8 Limitations and Future Work
Acknowledgements
Support for this project was provided by Google Research, PATH, and the Bill & Melinda Gates Foundation [Gates INV-068056]. We appreciate the excellent execution by the Intron Health Data Team who managed this pan-African data collection effort including recruiting contributors, developing custom UIs, quality reviews, and multi-currency contributor payments. We appreciate the invaluable support provided by the BioRAMP researchers, whose collaboration and insights have been fundamental to our research. We also extend our thanks to the Masakhane Community for their expert contributions, the Sisonke Biotik community for their enthusiastic participation and commitment, and the Federation of African Medical Students’ Associations [FAMSA] for their relentless recruitment and community engagement. These partnerships have been instrumental in enhancing the scope and impact of our work, and we appreciate their dedication to advancing knowledge in our field.

### v3

@@ Abstract
Recent advancements in large language model (LLM) performance on medical multiple-choice question (MCQ) benchmarks have stimulated interest from healthcare providers and patients globally. Particularly in low-and-middle-income countries (LMICs) facing acute physician shortages and lack of specialists, LLMs offer a potentially scalable pathway to enhance healthcare access and reduce costs. However, their effectiveness in the Global South, especially across the African continent, remains to be established. In this work, we introduce AfriMed-QA, the first large-scale Pan-African English multi-specialty medical Question-Answering (QA) dataset, 15,000 questions (open and closed-ended) sourced from over 60 medical schools across 16 countries, covering 32 medical specialties. We further evaluate 30 LLMs across multiple axes including correctness and demographic bias. Our findings show significant performance variation across specialties and geographies, MCQ performance clearly lags USMLE (MedQA). We find that biomedical LLMs underperform general models and smaller edge-friendly LLMs struggle to achieve a passing score. Interestingly, human evaluations show a consistent consumer preference for LLM answers and explanations when compared with clinician answers.
@@ 1 Introduction
To address this gap, we introduce AfriMed-QA, 22footnotetext: Available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias. We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches. While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs. The dataset is released under a CC-BY-NC-SA 4.0 license.
@@ 2.2 Evaluating LLMs for Health-Specific Tasks
Reddy (2023) proposed the TEHAI framework to assess the translational and governance aspects of LLMs in healthcare, emphasizing contextual relevance, safety, ethical considerations, and efficiency.
Our evaluation extends TEHAI and work from Singhal et al. (2022b), and Pfohl et al. (2024) by incorporating specialty and region-specific dimensions, ensuring a comprehensive assessment of LLMs. Metrics such as local expertise, harmfulness, and bias are included to cover ethical dimensions and governance. By addressing correctness, omission, hallucination, and reasonableness, we thoroughly evaluate AI systems in healthcare.
@@ 4.2 Quantitative Evaluation
Figure 3: AfriMedQA: Expert-MCQ accuracy, SAQ, and CQ Bertscore, sorted by Expert-MCQ Accuracy